OpenStudy (anonymous):

In the equation (1/3)ˣ = 2, the value of x rounded to the nearest thousandth is?

OpenStudy (lalaly):

\[\ln(\frac{1}{3})^x=\ln(2)\]\[xln(\frac{1}{3})=\ln(2)\]\[x=\frac{\ln(2)}{\ln(\frac{1}{3})}\]

OpenStudy (lalaly):

or u can say use \[b^x=a\]so\[\log_ba=x\]

OpenStudy (anonymous):

so in log form it would be xlog(1/3) = log(2)? Then dividing log(2) by log(1/3)?

OpenStudy (anonymous):

Okay yes the check works, thank you!

OpenStudy (anonymous):

take log on both sides